OpenStudy (anonymous):

Find the slope, if it exists 3x+5y=15

OpenStudy (saifoo.khan):

\[y = \frac{-3x+15}{5}\]

OpenStudy (saifoo.khan):

So, \[y = -\frac35x + 3\]

OpenStudy (saifoo.khan):

Slope = -3/5

OpenStudy (anonymous):

slope = -3/5

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

which equals -0.6

OpenStudy (saifoo.khan):

really?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thank you guys :)

OpenStudy (saifoo.khan):

Shawna is busy copying the answers/

OpenStudy (anonymous):

does anyone know the y-int and x-int of -3x+y=7

OpenStudy (saifoo.khan):

Yes i know.

OpenStudy (saifoo.khan):

Wanna know?

OpenStudy (anonymous):

yes please i been stuck on that one and 2 other ones for awhile

OpenStudy (saifoo.khan):

-3x+y=7 y = 3x + 7 Slope = 3 y-intercept = 7

OpenStudy (anonymous):

hey not nice

OpenStudy (saifoo.khan):

I was Jk.

OpenStudy (saifoo.khan):

Chill there!!

OpenStudy (anonymous):

meany

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

\[\Huge \text{i was jk.}\]

OpenStudy (anonymous):

i got two more problems and i will be done.... itz graphing though

OpenStudy (saifoo.khan):

Post them

OpenStudy (anonymous):

okay graph the equation 5x-4y=-8 and identify the y-int

OpenStudy (saifoo.khan):

SHanna knows the answers, she will help! ;)

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

y=2 slope = 5/2 which equals